OpenStudy (anonymous):

I'm thinking of a number. My number is 5 more than one-half my number. What is my number?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

your number (n) = 5 + (1/2)*n 2n = 10 + n n = 10

OpenStudy (mayankdevnani):

10 is correct.